OpenStudy (anonymous):

A(b,b) is sqrt(18) units from its origin. find b

OpenStudy (slaaibak):

This means sqrt(b^2 + b^2) = sqrt 18

OpenStudy (slaaibak):

Leading to \[2b^2 = 18\]

OpenStudy (slaaibak):

\[Distance = \sqrt{(x_2 - x_1)^2 + (y_2 - y_1)^2} \] where \[(b,b) = (x_2,y_2) \] and \[(0,0) = (x_1,y_1)\]

OpenStudy (amistre64):

since b,b is a 45 degree triangle with a hyp of sqrt(18).... id say b=sqrt(9)=3

OpenStudy (amistre64):

give or take some +- signs